# Contrastive Learning on Multimodal Analysis of Electronic Health Records

### v1

@@ 2.3 CLAIME Algorithm
We define code and CUI embedding matrices as $\mathbf{V}_{1}^{\star}=(\mathbf{v}_{1}^{\star},\ldots,\mathbf{v}_{d_{1}}^{\star})^{\top}\in\mathbb{R}^{d_{1}\times p}$ and $\mathbf{V}_{2}^{\star}=(\mathbf{v}_{d_{1}+1}^{\star},\ldots,\mathbf{v}_{d}^{\star})^{\top}\in\mathbb{R}^{d_{2}\times p}$, respectively, and aim to infer $\mathbf{V}^{\star}=\big{(}\mathbf{V}_{1}^{\star\top},\mathbf{V}_{2}^{\star\top}\big{)}^{\top}\in\mathbb{R}^{d\times p}$. The embeddings should reflect clinical semantics, meaning that highly similar (e.g. rheumatoid arthritis and juvenile rheumatoid arthritis) or related (e.g. fasting glucose and type II diabetes) EHR entities should have close embeddings. Before introducing our algorithm, we first define aggregate co-occurrence matrices $\mathbf{C}^{(M,M)}$ and $\mathbf{D}^{(M,M^{\prime})}$ for $M,M^{\prime}\in\{1,2\}$ across different modalities as:
@@
where $\mathbf{C}^{(M,M)}_{i,j}(w,w^{\prime})=\big{|}\{(t,s)\in[T_{i}^{(M)}]\times[T_{j}^{(M)}]:t\neq s,w^{(M)}_{i,t}=w,w^{(M^{\prime})}_{j,s}=w^{\prime}\}\big{|}$ and $\mathbf{D}^{(M,M^{\prime})}_{i,j}(w,w^{\prime})=\big{|}\{(t,s)\in[T_{i}^{(M)}]\times[T_{j}^{(M^{\prime})}]:w^{(M)}_{i,t}=w,w^{(M^{\prime})}_{j,s}=w^{\prime}\}\big{|}$ for $i,j\in[n]$ and $M\in\{1,2\}$. Further, we define the marginal co-occurrence of $w\in{\cal W}^{(M)}$ as:
@@
We introduce $S_{q}^{(M)}=(n^{-1}\sum_{i=1}^{n}(T_{i}^{(M)})^{q})^{1/q}$ for $q\geq 1$ and $S_{1}^{(1,2)}=n^{-1}\sum_{i=1}^{n}T_{i}^{(1)}T_{i}^{(2)}$ as scaling factors. CLAIME utilizes the multimodal contrastive learning loss defined as:
$\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})=\frac{1}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{1}^{(1,2)})}\sum_{i\neq j}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{j}^{(2)}]}\frac{\langle\mathbf{v}_{w^{(1)}_{i,t}},\mathbf{v}_{w^{(2)}_{j,s}}\rangle}{\gamma_{w_{i,t}}^{(1)}\gamma_{w_{j,s}}^{(2)}}$ (2.3)
$\quad\quad\quad\quad\quad\quad\quad\quad\quad-\frac{1}{nS_{1}^{(1,2)}}\sum_{i=1}^{n}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{i}^{(2)}]}\frac{\langle\mathbf{v}_{w^{(1)}_{i,t}},\mathbf{v}_{w^{(2)}_{i,s}}\rangle}{\gamma_{w_{i,t}}^{(1)}\gamma_{w_{i,s}}^{(2)}}+\frac{\lambda}{2}\|\mathbf{V}_{1}\mathbf{V}_{2}^{\top}\|_{\mathrm{F}}^{2}.$
Here $\mathbf{V}_{1}=(\mathbf{v}_{1},\ldots,\mathbf{v}_{d_{1}})^{\top}\in\mathbb{R}^{d_{1}\times p}$ and $\mathbf{V}_{2}=(\mathbf{v}_{d_{1}+1},\ldots,\mathbf{v}_{d})^{\top}\in\mathbb{R}^{d_{2}\times p}$, $\lambda>0$ serves as a regularization coefficient, $\gamma_{w}^{(M)},M=1,2$ are weights chosen based on the frequency of $w$, as defined in (2.2). Our theoretical analysis in Section 3 motivates the choice of $\gamma_{w}^{(M)}$ to guide the minimizer towards $\mathbf{V}^{\star}$. The essence of the multimodal contrastive learning loss in CLAIME is to enhance the representation of similar features across different modalities by bringing them closer together while distancing those that are dissimilar. In the context of EHR data, this translates to aligning the embeddings of codes and CUIs that have clinical correlations and separating those that do not. The aim is to maximize the inner product of embeddings for features that co-occur within the same patient’s data.
Remark 2.1.
Our CLAIME framework can be naturally extended to non-lienar loss functions. Note that our CLAIME loss function (2.3) can be written as
$\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ $=-\frac{1}{\sum_{i=1}^{n}T_{i}^{(1)}T_{i}^{(2)}}\sum_{i=1}^{n}\quantity{T_{i}^{(1)}T_{i}^{(2)}s_{ii}-\alpha_{T}\sum_{j:j\neq i}^{n}T_{i}^{(1)}T_{j}^{(2)}s_{ij}}+R(\mathbf{V}_{1},\mathbf{V}_{2}),$
where $R(\mathbf{V}_{1},\mathbf{V}_{2})$ is a smooth regularizer, and
$\alpha_{T}=\frac{\sum_{i=1}^{n}T_{i}^{(1)}T_{i}^{(2)}}{\sum_{i\neq j}^{n}T_{i}^{(1)}T_{j}^{(2)}},\ \ s_{ij}=\frac{1}{T_{i}^{(1)}T_{j}^{(2)}}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{j}^{(2)}]}\frac{\langle\mathbf{v}_{w^{(1)}_{i,t}},\mathbf{v}_{w^{(2)}_{j,s}}\rangle}{\gamma_{w_{i,t}}^{(1)}\gamma_{w_{j,s}}^{(2)}},\ \ \ \ \ \ i,j\in[n].$
One can consider the following non-linear loss function analogous to the CLIP loss function (Radford et al.,, 2021):
@@
$\mathcal{L}_{\textnormal{CLAIME}}^{\prime}(\mathbf{V}_{1},\mathbf{V}_{2})$ $=-\frac{1}{\sum_{i=1}^{n}T_{i}^{(1)}T_{i}^{(2)}}\sum_{i=1}^{n}T_{i}^{(1)}T_{i}^{(2)}\log\frac{\exp(s_{ii}/\tau)}{\sum_{j:j\neq i}^{n}T_{i}^{(1)}T_{j}^{(2)}\exp(s_{ij}/\tau)}+R(\mathbf{V}_{1},\mathbf{V}_{2}),$
which becomes equivalent to the loss (2.3) when $\tau\to\infty$.
To obtain $\mathbf{\widehat{V}}=(\mathbf{\widehat{V}}_{1}^{\top},\mathbf{\widehat{V}}_{2}^{\top})^{\top}=\arg\min\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ efficiently in practice, we note that $\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ can be expressed in terms of pair-wise co-occurrences of concepts as:
@@
$\frac{1}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{1}^{(1,2)})}\sum_{w=1}^{d_{1}}\sum_{w^{\prime}=d_{1}+1}^{d_{2}}\frac{\langle\mathbf{v}_{w},\mathbf{v}_{w^{\prime}}^{\star}\rangle}{\gamma_{w}^{(1)}\gamma_{w^{\prime}}^{(2)}}\sum_{i=1}^{n}\sum_{j\neq i}^{n}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{j}^{(2)}]}\mathbb{I}(w^{(1)}_{i,t}=w)\mathbb{I}(w^{(2)}_{j,s}=w^{\prime})$ (2.4)
$\quad-\frac{1}{nS_{1}^{(1,2)}}\sum_{w=1}^{d_{1}}\sum_{w^{\prime}=d_{1}+1}^{d_{2}}\frac{\langle\mathbf{v}_{w},\mathbf{v}_{w^{\prime}}^{\star}\rangle}{\gamma_{w}^{(1)}\gamma_{w^{\prime}}^{(2)}}\sum_{i=1}^{n}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{i}^{(2)}]}\mathbb{I}(w^{(1)}_{i,t}=w)\mathbb{I}(w^{(2)}_{i,s}=w^{\prime})+\frac{\lambda}{2}\|\mathbf{V}_{1}\mathbf{V}_{2}^{\top}\|_{\mathrm{F}}^{2}\,.$
Subsequently, via arguments given in Supplementary S2, we have the following proposition.
@@
$\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ $=\frac{\lambda}{2}\Big{\|}\mathbf{V}_{1}\mathbf{V}_{2}^{\top}-\frac{1}{\lambda}\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}\Big{\|}_{\mathrm{F}}^{2}+(\textnormal{constant}),$
@@
$\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}(w,w^{\prime})$ $:=\frac{\mathbf{C}^{(1,1)}(\cdot,\cdot)\mathbf{C}^{(2,2)}(\cdot,\cdot)}{\mathbf{C}^{(1,1)}(w,\cdot)\mathbf{C}^{(2,2)}(w^{\prime},\cdot)}\quantity(\frac{\mathbf{D}^{(1,2)}(w,w^{\prime})}{\mathbf{D}^{(1,2)}(\cdot,\cdot)}-\frac{\mathbf{C}^{(c)}(w,w^{\prime})}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{1}^{(1,2)})}),$
and $\mathbf{C}^{(c)}(w,w^{\prime})=\sum_{i=1}^{n}\sum_{j\neq i}^{n}\mathbf{D}^{(1,2)}_{i,j}(w,w^{\prime}).$
Proposition 2.1 shows that $\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ can be related to the SVD of an empirical association matrix, where its element $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}(w,w^{\prime})$ estimates the association between the features $w$ and $w^{\prime}$. We will later demonstrate the convergence of $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$ to the population PMI matrix in Section 3. As a result, $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}(w,w^{\prime})$ can be viewed as a modified estimator of the population PMI matrix. The final embeddings $\mathbf{\widehat{V}}_{1}\mathbf{\widehat{V}}_{2}^{\top}$ are inferred through a rank-$p$ SVD of $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$, preserving data privacy and offering a scalable estimation. To be more specific, let $\widehat{\mathbf{U}}_{1}\widehat{\bm{\Lambda}}\widehat{\mathbf{U}}_{2}$ denote the rank-$p$ SVD of $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$, where $\widehat{\mathbf{U}}_{1}\in\mathbb{R}^{d_{1}\times p}$ and $\widehat{\mathbf{U}}_{2}\in\mathbb{R}^{d_{2}\times p}$ are the matrices of left and right singular vectors, respectively, and $\widehat{\bm{\Lambda}}\in\mathbb{R}^{p\times p}$ is a diagonal matrix with its diagonal elements being the top $p$ singular values. Then, we set $\mathbf{\widehat{V}}_{1}=\widehat{\mathbf{U}}_{1}\widehat{\bm{\Lambda}}^{1/2}$ and $\mathbf{\widehat{V}}_{2}=\widehat{\mathbf{U}}_{2}\widehat{\bm{\Lambda}}^{1/2}$. It is worth noting that the selection of the regularization parameter $\lambda$ does not play a crucial role, and for the sake of simplicity, we will assign it a value of $1$ in our upcoming numerical analyses.
2.4 Comparison between CLAIME and Simple Concatenation

### v2

@@ 2.2 The CLAIME Algorithm
We define code and CUI embedding matrices as $\mathbf{V}_{1}^{\star}=(\mathbf{v}_{1}^{\star},\ldots,\mathbf{v}_{d_{1}}^{\star})^{\scriptscriptstyle\sf T}\in\mathbb{R}^{d_{1}\times p}$ and $\mathbf{V}_{2}^{\star}=(\mathbf{v}_{d_{1}+1}^{\star},\ldots,\mathbf{v}_{d_{1}+d_{2}}^{\star})^{\scriptscriptstyle\sf T}\in\mathbb{R}^{d_{2}\times p}$, respectively, and aim to infer $\mathbf{V}^{\star}=\big{(}\mathbf{V}_{1}^{\star{\scriptscriptstyle\sf T}},\mathbf{V}_{2}^{\star{\scriptscriptstyle\sf T}}\big{)}^{\scriptscriptstyle\sf T}\in\mathbb{R}^{d\times p}$. The embeddings are intended to reflect clinical semantics, meaning that highly similar (e.g., rheumatoid arthritis and juvenile rheumatoid arthritis) or related (e.g., fasting glucose and type II diabetes) EHR entities should be embedded close to each other.
To motivate our algorithm design, recall that model (2.1) assumes EHR features from both modalities are generated from a shared latent patient embedding $\bm{c}_{i}$. This implies that features frequently co-occurring within the same patient are likely conditionally dependent through $\bm{c}_{i}$, and should have similar embeddings, whereas features from different patients are marginally independent and their embeddings should remain unaligned.
Before introducing our algorithm, we first define aggregate co-occurrence matrices $\mathbf{C}^{(M,M)}$ and $\mathbf{D}^{(M,M^{\prime})}$ for $M,M^{\prime}\in\{1,2\}$ across different modalities as:
@@
where $\mathbf{C}^{(M,M)}_{i,j}(w,w^{\prime})=\big{|}\{(t,s)\in[T_{i}^{(M)}]\times[T_{j}^{(M)}]:t\neq s,w^{(M)}_{i,t}=w,w^{(M^{\prime})}_{j,s}=w^{\prime}\}\big{|}$ and $\mathbf{D}^{(M,M^{\prime})}_{i,j}(w,w^{\prime})=\big{|}\{(t,s)\in[T_{i}^{(M)}]\times[T_{j}^{(M^{\prime})}]:w^{(M)}_{i,t}=w,w^{(M^{\prime})}_{j,s}=w^{\prime}\}\big{|}$ for $i,j\in[n]$. Further, we define the marginal occurrence count of $w\in{\cal W}^{(M)}$ as:
@@
To derive clinically meaningful embeddings that reflect modality-specific co-occurrence patterns, we propose a contrastive learning algorithm aligned with model (2.1). Contrastive learning encourages embeddings of positive pairs to be similar and those of negative pairs to be dissimilar. In our setting, code-CUI pairs from the same patient are treated as positive, while pairs from different patients as negative. This naturally aligns with our generative assumptions and enables learning from aggregated co-occurrence data. Specifically, we define the multimodal contrastive learning loss as
@@
$\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})=\frac{\sum_{i,j:i\neq j}s_{ij}}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{2}^{(1,2)2})}-\frac{\sum_{i=1}^{n}s_{ii}}{nS_{2}^{(1,2)2}}+\frac{\lambda}{2}\|\mathbf{V}_{1}\mathbf{V}_{2}^{\scriptscriptstyle\sf T}\|_{\mathrm{F}}^{2},$ (2.3)
where $s_{ij}=\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{j}^{(2)}]}\frac{\langle\mathbf{v}_{w^{(1)}_{i,t}},\mathbf{v}_{w^{(2)}_{j,s}}\rangle}{\gamma_{w_{i,t}^{(1)}}^{(1)}\gamma_{w_{j,s}^{(2)}}^{(2)}}$. Here, $\mathbf{V}_{1}=(\mathbf{v}_{1},\ldots,\mathbf{v}_{d_{1}})^{\scriptscriptstyle\sf T}\in\mathbb{R}^{d_{1}\times p}$ and $\mathbf{V}_{2}=(\mathbf{v}_{d_{1}+1},\ldots,\mathbf{v}_{d_{1}+d_{2}})^{\scriptscriptstyle\sf T}\in\mathbb{R}^{d_{2}\times p}$, $\lambda>0$ serves as a regularization coefficient, $\gamma_{w}^{(M)},M=1,2$ are weights chosen based on the frequency of $w$, as defined in (2.2). The quantities $S_{q}^{(M)}=(n^{-1}\sum_{i=1}^{n}(T_{i}^{(M)})^{q})^{1/q}$ and $S_{q}^{(1,2)}=(n^{-1}\sum_{i\in[n]}(T_{i}^{(1)})^{q/2}(T_{i}^{(2)})^{q/2})^{1/q}$ for $q\geq 1$ are empirical moments and cross-moments of $T_{i}^{(M)}$, and correspond to the number of terms in the two summation components of the loss. Our analysis in Section 3 motivates the choice of $\gamma_{w}^{(M)}$ to guide the minimizer towards $\mathbf{V}^{\star}$.
This loss encourages code-CUI pairs that co-occur within the same patient (second term) to be closely embedded, while penalizing similarity between pairs from different patients (the first term). The regularization term $\frac{\lambda}{2}\|\mathbf{V}_{1}\mathbf{V}_{2}^{\scriptscriptstyle\sf T}\|_{\mathrm{F}}^{2}$ ensures numerical stability and controls the embedding magnitude. By optimizing this loss, the model learns to align embeddings based on co-occurrence patterns, which signal shared clinical relevance.
To efficiently obtain $\mathbf{\widehat{V}}=(\mathbf{\widehat{V}}_{1}^{\scriptscriptstyle\sf T},\mathbf{\widehat{V}}_{2}^{\scriptscriptstyle\sf T})^{\scriptscriptstyle\sf T}=\arg\min\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$, we define normalized count vectors for each patient $i$ and modality $M\in\{1,2\}$: $\mathbf{w}_{i}^{(M)}\in\mathbb{R}^{d_{M}}$ with $\mathbf{w}_{i,w}^{(M)}=\frac{1}{\gamma_{w}^{(M)}}\sum_{t\in[T_{i}^{(M)}]}\mathbb{I}(w_{i,t}^{(M)}=w)$. The patient-level representation for each modality becomes $\mathbf{V}_{1}^{\scriptscriptstyle\sf T}\mathbf{w}_{i}^{(1)}$ and $\mathbf{V}_{2}^{\scriptscriptstyle\sf T}\mathbf{w}_{i}^{(2)}\in\mathbb{R}^{p}$. Using this, the loss (2.3) (excluding regularization) can be expressed by pair-wise co-occurrences of concepts as:
$\frac{1}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{2}^{(1,2)2})}\sum_{w=1}^{d_{1}}\sum_{w^{\prime}=d_{1}+1}^{d_{1}+d_{2}}\frac{\langle\mathbf{v}_{w},\mathbf{v}_{w^{\prime}}\rangle}{\gamma_{w}^{(1)}\gamma_{w^{\prime}}^{(2)}}\sum_{i=1}^{n}\sum_{j:j\neq i}^{n}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{j}^{(2)}]}\mathbb{I}(w^{(1)}_{i,t}=w)\mathbb{I}(w^{(2)}_{j,s}=w^{\prime})$
@@
$\quad-\frac{1}{nS_{2}^{(1,2)2}}\sum_{w=1}^{d_{1}}\sum_{w^{\prime}=d_{1}+1}^{d_{1}+d_{2}}\frac{\langle\mathbf{v}_{w},\mathbf{v}_{w^{\prime}}\rangle}{\gamma_{w}^{(1)}\gamma_{w^{\prime}}^{(2)}}\sum_{i=1}^{n}\sum_{t\in[T_{i}^{(1)}]}\sum_{s\in[T_{i}^{(2)}]}\mathbb{I}(w^{(1)}_{i,t}=w)\mathbb{I}(w^{(2)}_{i,s}=w^{\prime})$
$=\frac{1}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{2}^{(1,2)2})}\sum_{i=1}^{n}\sum_{j:j\neq i}^{n}(\mathbf{V}_{1}^{\scriptscriptstyle\sf T}\mathbf{w}_{i}^{(1)})^{\scriptscriptstyle\sf T}(\mathbf{V}_{2}^{\scriptscriptstyle\sf T}\mathbf{w}_{j}^{(2)})-\frac{1}{nS_{2}^{(1,2)2}}\sum_{i=1}^{n}(\mathbf{V}_{1}^{\scriptscriptstyle\sf T}\mathbf{w}_{i}^{(1)})^{\scriptscriptstyle\sf T}(\mathbf{V}_{2}^{\scriptscriptstyle\sf T}\mathbf{w}_{i}^{(2)}).$
The first term captures the average similarity of negative (cross-patient) pairs, while the second reflects positive (same-patient) pairs. Minimizing this loss encourages within-patient alignment while discouraging cross-patient alignment. Subsequently, via arguments given in Supplementary S2, we have the following proposition.
@@
$\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ $=\frac{\lambda}{2}\Big{\|}\mathbf{V}_{1}\mathbf{V}_{2}^{\scriptscriptstyle\sf T}-\frac{1}{\lambda}\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}\Big{\|}_{\mathrm{F}}^{2}+(\textnormal{constant}),$
@@
$\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}(w,w^{\prime}):=\frac{\mathbf{C}^{(1,1)}(\cdot,\cdot)\mathbf{C}^{(2,2)}(\cdot,\cdot)}{\mathbf{C}^{(1,1)}(w,\cdot)\mathbf{C}^{(2,2)}(w^{\prime},\cdot)}\quantity(\frac{\mathbf{D}^{(1,2)}(w,w^{\prime})}{\mathbf{D}^{(1,2)}(\cdot,\cdot)}-\frac{\mathbf{C}^{(c)}(w,w^{\prime})}{n(nS_{1}^{(1)}S_{1}^{(2)}-S_{2}^{(1,2)2})}),$
and $\mathbf{C}^{(c)}(w,w^{\prime})=\sum_{i=1}^{n}\sum_{j:j\neq i}^{n}\mathbf{D}^{(1,2)}_{i,j}(w,w^{\prime}).$
Proposition 2.1 shows that $\mathcal{L}_{\textnormal{CLAIME}}(\mathbf{V}_{1},\mathbf{V}_{2})$ reduces to a rank-constrained approximation of an empirical association matrix $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$, where each entry $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}(w,w^{\prime})$ quantifies the association between features $w$ and $w^{\prime}$. Section 3 establishes that this matrix converges to the population-level PMI matrix under mild assumptions.
To estimate the embeddings, we compute the rank-$p$ SVD of $\widehat{\mathbb{PMI}}_{\textnormal{CLAIME}}$, denoted as $\widehat{\mathbf{U}}_{1}\widehat{\bm{\Lambda}}\widehat{\mathbf{U}}_{2}^{{\scriptscriptstyle\sf T}}$, where $\widehat{\mathbf{U}}_{1}\in\mathbb{R}^{d_{1}\times p}$ and $\widehat{\mathbf{U}}_{2}\in\mathbb{R}^{d_{2}\times p}$ are the matrices of left and right singular vectors, respectively, and $\widehat{\bm{\Lambda}}\in\mathbb{R}^{p\times p}$ is a diagonal matrix with its diagonal elements being the top $p$ singular values. Then, we set $\mathbf{\widehat{V}}_{1}=\widehat{\mathbf{U}}_{1}\widehat{\bm{\Lambda}}^{1/2}$ and $\mathbf{\widehat{V}}_{2}=\widehat{\mathbf{U}}_{2}\widehat{\bm{\Lambda}}^{1/2}$. From Proposition 2.1, the scalar $\lambda$ in (2.3) controls the scale of the embeddings via the magnitude of singular values, but not their subspace. Since downstream similarity metrics like cosine similarity are scale-invariant, we may fix $\lambda=1$ for simplicity.
Our formulation is inspired by contrastive learning, which promotes high similarity among positive pairs and low similarity among negatives. Unlike vision-language models such as CLIP (Radford et al., , 2021), it operates on discrete, high-dimensional EHR features rather than continuous image or text embeddings, and relies only on aggregated co-occurrence statistics instead of raw patient-level data.
Remark 2.2 (Nonlinear extension).
The CLAIME framework can be naturally extended to a softmax-based contrastive loss similar to that used in CLIP:
$\mathcal{L}_{\textnormal{CLAIME}}^{\prime}(\mathbf{V}_{1},\mathbf{V}_{2})$ $=-\frac{1}{n(S_{2}^{(1,2)})^{2}}\sum_{i=1}^{n}T_{i}^{(1)}T_{i}^{(2)}\log\frac{\exp(s_{ii}/(T_{i}^{(1)}T_{i}^{(2)}\eta))}{\sum_{j:j\neq i}T_{i}^{(1)}T_{j}^{(2)}\exp(\alpha_{i}s_{ij}/(T_{i}^{(1)}T_{j}^{(2)}\eta))}+R^{\prime}(\mathbf{V}_{1},\mathbf{V}_{2}),$ (2.4)
where $R^{\prime}$ is a smooth regularizer, $\eta>0$ the temperature, and $\alpha_{i}$ depends on $(T_{i}^{(M)})_{i\in[n],M=1,2}$. In the high-temperature limit ($\eta\to\infty$), the loss reduces to the linear CLAIME loss (2.3), up to a constant and regularization. Details are provided in Supplementary S2.2.
Remark 2.3 (Extending CLAIME to more than two modalities).
EHR data often consist of multiple heterogeneous streams. For instance, codified data alone encompass diagnoses, procedures, medications, and laboratory results, each of which can be treated as a distinct modality. Similarly, NLP-derived CUIs can be partitioned by semantic types such as Disease or Syndrome and Diagnostic Procedure. To accommodate such settings, we generalize CLAIME beyond two modalities in Supplementary S6 by estimating cross-modal PMI matrices for every pair of modalities and jointly learning embeddings over all sources. Further extensions to continuous-valued sources such as genomics, imaging, or clinical tests will require suitable generative models for continuous data.
2.3 Comparison between CLAIME and Baseline Approaches